# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

### v1

@@ 4 Experiments
Training Data & Setting. We develop HotpotQA-SUBQ, a dataset derived from HotpotQA (Yang et al., 2018), to train our model’s action planning and answering capabilities. Our dataset creation begins with document filtering: using Claude 3.5 Sonnet (Anthropic, 2024), we identify and retain only the supporting documents necessary for answering the main query, removing irrelevant content. For each supporting document , we then iteratively generate a subquery , considering the main query, previous subqueries, and supporting documents. During iteration , when more supporting documents remain, we create training samples with input and output label “[SUBQ] ”, where represents triples extracted from document , and is the main query. For the final supporting document, we label the sample as “[SUFFICIENT]”. To identify queries that can be answered directly, we test our base LLM (LLaMA-2-7B) on HotpotQA’s main queries without context. For correctly answered queries, we create training samples with the main query as input and “[NO_RETRIEVAL]” as the output label. Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from SelfRAG’s training data, resulting in 208k training samples in total. We place detailed training data processing, dataset statistics, and data samples in Appendix B. We use DBPedia-298, HotpotQA-SUBQ, and WikiOFGraph to train our theme classifier , distribution shifter , and text-to-triples model , respectively, as detailed in Appendix C. We present our hyperparameter study of each component in Appendix G.
@@
Knowledge Sources. Due to the large sizes of wiki corpora, we employ faiss (Douze et al., 2024) for efficient vector searching across both theme and dense indices. Following the SelfRAG (Asai et al., 2023), we utilize the Wikipedia 2018 (Izacard et al., 2023) by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information. To optimize retrieval efficiency, we partition the index into five segments. Within each segment, we first perform theme scoping to reduce the candidate document pool to 250,000 documents by default. However, we skip theme scoping for PopQA since our classifier , trained on DBPedia-298 data based on old Wiki dump, performs poorly on newer Wiki 2020 dump, producing near-zero vectors for roughly one-third of recent content.
@@
Test Datasets & Metrics & Compared Baselines. We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies (Asai et al., 2023; Lyu et al., 2024b). The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2022), and 2WikiMultihopQA (Ho et al., 2020); (2) closed-set task datasets: PubHealth (Zhang et al., 2023a) and ARC-Challenge (Clark et al., 2018); and (3) long-form generation datasets: ALCE-ASQA (Gao et al., 2023; Stelmakh et al., 2022) and ELI5 (Fan et al., 2019). For evaluation metrics, we maintain consistency with prior work (Asai et al., 2023; Mallen et al., 2022; Lyu et al., 2024b), employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score (Pillutla et al., 2021) for ASQA and ELI5. Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe (Kim et al., 2024b), a leading retrieve-and-summarize framework; and models with self-reflective retrieval, including leading approaches Self-RAG (Asai et al., 2023) and RPG (Lyu et al., 2024b). We place more details of metrics/datasets in Appendix D.
| Short-form | Closed-set | Long-form Generation | ||||||||
| Model | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | |||
| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | ||
| Baselines without retrieval | ||||||||||
| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 | |
| Alpaca | 54.5 | – | 23.6 | 49.8 | 45.0 | 29.4 | 61.7 | – | – | |
| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 | |
| Alpaca | 61.3 | – | 24.4 | 55.5 | 54.9 | 32.0 | 70.6 | – | – | |
| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 | |
| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 | |
| Baselines with retrieval (#docs=5 by default) | ||||||||||
| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 | |
| Alpaca | 64.1 | – | 46.7 | 40.2 | 48.0 | 33.3 | 57.9 | – | – | |
| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 | |
| Alpaca | 66.9 | – | 46.1 | 51.1 | 57.6 | 36.7 | 56.6 | – | – | |
| RECOMP (Xu et al., 2023) | – | 32.4 | – | – | – | 36.5 | 76.0 | – | – | |
| SuRe (Kim et al., 2024b) | – | 20.6 | – | – | – | 35.8 | 76.2 | – | – | |
| SuRe | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | – | – | |
| ChatGPT | 65.7 | 32.8 | 50.8 | 54.7 | 75.3 | 39.9 | 79.7 | 20.6 | 57.2 | |
| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | – | – | – | – | |
| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 | |
| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 | |
| Baselines with self-reflective retrieval | ||||||||||
| SELF-RAG (Asai et al., 2023) | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 | |
| SELF-RAG | 69.3 | – | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | – | – | |
| RPG (Lyu et al., 2024b) | – | 33.6 | 56.0 | 73.4 | – | 37.6 | 84.4 | 19.1 | 46.4 | |
| Retrieval-And-Structuring (Ours) | ||||||||||
| RAS | RAS | 77.4 | 57.7 | 62.3 | 71.4 | 93.8 | 39.1 | 70.5 | 23.3 | 37.7 |
| RAS | RAS | 73.8 | 43.4 | 61.3 | 75.5 | 62.4 | 37.2 | 95.2 | 19.7 | 47.8 |
| RAS | RAS | 78.7 | 54.7 | 58.8 | 70.8 | 91.2 | 39.1 | 70.5 | 23.3 | 37.7 |
| RAS | RAS | 72.6 | 42.1 | 58.3 | 74.8 | 67.7 | 37.2 | 95.2 | 19.7 | 47.8 |
[…]
| RAS | RAS | 73.8 | 44.2 | 57.7 | 77.5 | 71.5 | 37.6 | 96.2 | 20.1 | 54.4 |
@@
Inference Setting. We evaluate RAS using both our trained open-source model (RAS with LLaMA-2-7B/LLaMA-3-8B and Graph Transformer (Shi et al., 2020) encoder) and the closed-source Claude-3.5-Sonnet model under varied inference settings. For RAS, we maintain consistency with previous work (Asai et al., 2023; Lyu et al., 2024b) by employing zero-shot inference across all datasets. For closed-source models (e.g., RAS), we utilize few-shot inference with two exemplars for ASQA and ELI5, while maintaining zero-shot for the other datasets.
@@
During inference, if the initial planner output is not [NO_RETRIEVAL], we utilize the main query as the initial subquery . For PopQA and TriviaQA evaluation, we follow established settings (Asai et al., 2023; Luo et al., 2023), incorporating top-five web search engine results as initial retrieved context . For ASQA and ELI5, we maintain methodological consistency with prior work (Lyu et al., 2024b; Asai et al., 2023; Gao et al., 2023), utilizing their predetermined five-document context. Under these conditions, we omit plan generation and text retrieval phases, implementing static inference (Asai et al., 2023) with five fixed iterations. For remaining datasets, we establish a maximum iteration count of five. Following previous studies, we employ Contriever-MS MARCO (Izacard et al., 2021) as the primary dense retriever, with BM25 (Robertson et al., 2009) serving as the retrieval mechanism for 2WikiMultihopQA. Across all retrieval processes, we maintain a consistent top- document selection of five. We present more details of inference settings in Appendix E.
@@
| No GraphToken | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
| No Planning | 66.7 | 37.8 | 71.5 | – | – |
| No ThemeScope | 72.0 | – | 73.9 | – | – |
4.1 Main Results
Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing open-source solutions, including SelfRAG (Asai et al., 2023) and RPG (Lyu et al., 2024b). Notably, compared to the previous SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks. Additionally, when applied to closed-source Sonnet-3.5 model, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe (Kim et al., 2024b).
We find that sometimes (e.g., on TriviaQA and PubHealth) single-hop retrieval could not boost LLM’s performance, and even makes it worse, which demonstrates the necessity of on-demand retrieval, aligning with previous findings.
Our “role-swapping” study, where Sonnet-3.5 and RAS alternately serve as action planner and answerer, reveals that performance is primarily bounded by the answerer’s inherent capabilities. For example, when Sonnet-3.5 acts as the planner with RAS as the answerer, the system achieves 62.4% accuracy on the ARC-Challenge dataset. However, when Sonnet-3.5 serves as both planner and answerer, accuracy increases to 93.8%. Notably, despite having 60x fewer parameters, our trained RAS performs comparably to Sonnet-3.5 as a planner, and even outperforms it sometimes.
Also, for long-form generation task, our trained model achieved notably high MAUVE scores, which can be majorly attributed to our “text-to-triple” generation step, as demonstrated by our later ablation study in Table 2.
4.2 Ablations & More Analysis
We study the effect of individual component in Table 2.
@@ 4.2 Ablations & More Analysis
Effect of LoRA and Multi-task Learning. Our experiments reveal that parameter-efficient training strategies significantly impact model performance. Using only graph token optimization without LoRA leads to substantial degradation (-11.8% on average). A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work (Lyu et al., 2024b). The complementary effects suggest that while graph-based knowledge representation is valuable, it needs to be combined with careful parameter tuning and multi-task learning to achieve optimal performance.
Effect of Theme Scoping. The theme-scoped retrieval shows mixed results – while removing it (“No ThemeScoping”) has minimal performance drops on TQA and PubHealth, the efficiency benefits are substantial as shown in Figure 2. This suggests theme scoping successfully narrows the search space without compromising performance.
Efficiency Analysis. Figure 2 compares the computational costs of different retrieval strategies. Theme-scoped RAG reduces processing time by approximately 60% compared to dense-only RAG by first filtering the corpus through lightweight theme classification (with 298 class dimensions). This combination of comparable performance with significantly reduced computational cost demonstrates the effectiveness of our theme-scoped retrieval approach.
@@
Impact of Training Data Volume. Figure 3 demonstrates how training dataset size influences model performance across different tasks. Considering the computational efficiency, we sampled 2,000 instances each from TQA and 2WQA for evaluation, while maintaining the original sizes of other datasets. The results indicate that model performance generally improves with increased training data volume. Notably, our model achieves competitive performance even with limited training data - using only 5% (10K instances) of our full dataset, it surpasses previous state-of-the-art models on TQA, 2WQA, and ELI5. These results suggest both the robustness of our architectural design and the effectiveness of our data curation methodology. However, we observed an exception with the ELI5 dataset, where performance were inconsistent. This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research (Asai et al., 2023; Lyu et al., 2024b). Among our test datasets, ASQA and ELI5 are unique in requiring long-form response generation. The periodic decline in ELI5 performance suggests that the model’s response generation began to align more closely with ASQA’s training data distribution, potentially at the expense of ELI5’s distinct characteristics.
Impact of Graph Information Abundance. Figure 4 demonstrates a strong positive correlation between graph information abundance and model performance across all tasks. Both RAS and RAS show consistent improvements as graph information increases, demonstrating RAS’s strong capability of effective utilizing available graph information for knowledge-intensive generation.

### v2

@@ 4 Experiments
Training Data & Setting. We develop HotpotQA-SUBQ, a dataset constructed based on HotpotQA [1], to train our model’s action planning and answering capabilities. Our dataset creation begins with document filtering: using Claude-3.5-Sonnet [48], we identify and retain only the supporting documents necessary for answering the main query, removing irrelevant content. For each supporting document , we then iteratively generate a subquery , considering the main query, previous subqueries, and supporting documents. During iteration , when more supporting documents remain, we create training samples with input and output label “[SUBQ] ”, where represents triples extracted from document , and is the main query. For the final supporting document, we label the sample as “[SUFFICIENT]”. To identify queries that can be answered directly, we test our base LLM (LLaMA-2-7B) on HotpotQA’s main queries without context. For correctly answered queries, we create training samples with the main query as input and “[NO_RETRIEVAL]” as the output label. Additionally, to ensure fair comparison with existing approaches, we incorporate the subset of Arc-Easy (2,147 samples) and ASQA (3,897 samples) from Self-RAG’s training data, resulting in 208k training samples in total. We place detailed training data processing, dataset statistics, and data samples in Appendix B. For efficient inference, we train a text-to-triples model on the WikiOFGraph dataset [44] using LLaMA-3.2-3B-Instruct as the base model (see Appendix C), and deploy it using vLLM for optimized runtime performance. We present our hyperparameter study of each component in Appendix H.
@@
Knowledge Sources. Due to the large sizes of wiki corpora, we employ faiss [49] for efficient vector searching over the dense index. Following the Self-RAG [30], we utilize the Wikipedia 2018 [50] by default, while specifically using the Wikipedia 2020 for PopQA to access more recent information. To optimize retrieval efficiency, we partition the index into five segments.
@@
Test Datasets & Metrics & Compared Baselines. We conduct comprehensive evaluations on diverse knowledge-intensive tasks following previous studies [30, 31]. The evaluation encompasses three categories of datasets: (1) open-domain short-form generation datasets: TriviaQA [51], PopQA [52], and 2WikiMultihopQA [53]; (2) closed-set task datasets: PubHealth [54] and ARC-Challenge [55]; and (3) long-form generation datasets: ALCE-ASQA [56, 57] and ELI5 [58]. For evaluation metrics, we maintain consistency with prior work [30, 52, 31], employing “golden match” accuracy for PopQA and TriviaQA, token-level F1 score for 2WikiMultihopQA, accuracy for PubHealth and ARC-Challenge, and ROUGE-LSum alongside MAUVE score [59] for ASQA and ELI5. Our comparative analysis includes three baseline categories: models without retrieval augmentation, incorporating Claude 3.5 Sonnet as a state-of-the-art closed-source baseline; models with single retrieval over top-5 documents, including Claude 3.5 Sonnet, and SuRe [25], a leading retrieve-and-summarize method; and models with self-reflective retrieval, including leading approaches Self-RAG [30], RPG [31], and Search-R1 [60]. We place more details of metrics/datasets in Appendix D.
| Short-form | Closed-set | Long-form Generation | |||||||
| Model | TQA | 2WQA | PopQA | Pub | ARC | ASQA | ELI5 | ||
| (acc) | (F1) | (acc) | (acc) | (acc) | (rouge) | (mauve) | (rouge) | (mauve) | |
| Baselines w/o Retrieval | |||||||||
| Llama2 | 30.5 | 18.9 | 14.7 | 34.2 | 21.8 | 15.3 | 19.0 | 18.3 | 32.4 |
| Alpaca | 54.5 | – | 23.6 | 49.8 | 45.0 | 29.4 | 61.7 | – | – |
| Llama2 | 38.5 | 20.2 | 14.7 | 29.4 | 29.4 | 12.4 | 16.0 | 18.2 | 41.4 |
| Alpaca | 61.3 | – | 24.4 | 55.5 | 54.9 | 32.0 | 70.6 | – | – |
| Llama3 | 56.1 | 21.2 | 26.7 | 33.2 | 42.2 | 17.6 | 25.0 | 18.2 | 39.7 |
| ChatGPT | 74.3 | 24.8 | 29.3 | 70.1 | 75.3 | 36.2 | 68.8 | 22.8 | 32.6 |
| Sonnet-3.5 | 78.4 | 40.0 | 30.2 | 83.7 | 88.5 | 37.0 | 39.1 | 21.8 | 26.5 |
| Baselines w/ Retrieval (#docs=5 by default) | |||||||||
| Llama2 | 42.5 | 21.0 | 38.2 | 30.0 | 48.0 | 22.1 | 32.0 | 18.6 | 35.3 |
| Alpaca | 64.1 | – | 46.7 | 40.2 | 48.0 | 33.3 | 57.9 | – | – |
| Llama2 | 47.0 | 31.2 | 45.7 | 30.2 | 26.0 | 20.5 | 24.7 | 18.6 | 42.3 |
| Llama3 | 60.4 | 33.4 | 48.6 | 36.5 | 40.1 | 23.9 | 52.1 | 18.8 | 40.7 |
| Alpaca | 66.9 | – | 46.1 | 51.1 | 57.6 | 36.7 | 56.6 | – | – |
| RECOMP [61] | – | 32.4 | – | – | – | 36.5 | 76.0 | – | – |
| ChatGPT | 65.7 | 32.8 | 50.8 | 54.7 | 75.3 | 39.9 | 79.7 | 20.6 | 57.2 |
| Sonnet-3.5 | 69.1 | 41.9 | 51.5 | 49.1 | 88.6 | – | – | – | – |
| Sonnet-3.5 | 72.5 | 53.7 | 57.3 | 53.9 | 87.1 | 38.8 | 61.6 | 20.2 | 32.3 |
| SuRe [25] | 51.2 | 20.6 | 39.0 | 36.2 | 52.7 | 35.8 | 76.2 | 16.1 | 26.6 |
| SuRe | 72.3 | 38.1 | 53.6 | 57.2 | 79.6 | 36.0 | 74.2 | 19.2 | 51.6 |
| SuRe | 76.8 | 37.6 | 41.2 | 62.8 | 91.6 | 30.2 | 69.9 | 15.4 | 27.2 |
| Baselines w/ Self-Reflective Retrieval | |||||||||
| Self-RAG [30] | 66.4 | 25.1 | 54.9 | 72.4 | 67.3 | 35.7 | 74.3 | 17.9 | 35.6 |
| Self-RAG | 69.3 | – | 55.8 | 74.5 | 73.1 | 37.0 | 71.6 | – | – |
| RPG [31] | 65.1 | 33.6 | 56.0 | 73.4 | 65.4 | 37.6 | 84.4 | 19.1 | 46.4 |
| Search-R1 [60] | 63.8 | 35.2 | 45.7 | 69.4 | 64.4 | 35.2 | 72.3 | 18.4 | 37.1 |
| Retrieval-And-Structuring (RAS) | |||||||||
| RAS | 77.6 | 57.7 | 62.3 | 71.3 | 93.9 | 39.1 | 70.5 | 23.3 | 37.7 |
| RAS | 73.8 | 44.2 | 57.7 | 77.6 | 71.4 | 37.6 | 96.2 | 20.1 | 54.4 |
| RAS | 72.7 | 42.1 | 58.3 | 74.7 | 68.5 | 37.2 | 95.2 | 19.7 | 47.8 |
@@
Inference Setting. We evaluate RAS using both our trained open-source models (RAS, based on LLaMA-2-7B/LLaMA-3-8B and a Graph Transformer encoder [65]) and the closed-source Claude-3.5-Sonnet model under varied inference settings. For open-source models (RAS), we follow prior work [30, 31] and adopt zero-shot inference across all datasets. For the closed-source model (RAS), we apply few-shot inference with two exemplars for ASQA and ELI5, while using zero-shot inference for all other datasets. More details are provided in Appendix E.1 and E.2.
@@
During inference, if the initial planning output is not [NO_RETRIEVAL], we utilize the main query as the initial subquery . For PopQA and TriviaQA evaluation, we follow established settings [30, 66], incorporating top-five web search engine results as initial retrieved context . For ASQA and ELI5, we maintain methodological consistency with prior work [31, 30, 56], utilizing their predetermined five-document context. Under these conditions, we omit plan generation and text retrieval phases, implementing static inference [30] with five fixed iterations. For remaining datasets, we establish a maximum iteration count of five. Following previous studies, we employ Contriever-MS MARCO [67] as the primary dense retriever, with BM25 [68] serving as the retrieval mechanism for 2WikiMultihopQA. Across all retrieval processes, we maintain a consistent top- document selection of five. We present more details of inference settings in Appendix E.
4.1 Main Results
Our performance evaluation, presented in Table 1, demonstrates that the LLaMA-2-7B/LLaMA-3-8B model fine-tuned with RAS outperforms existing open-source solutions, including Self-RAG [30] and RPG [31]. Notably, compared to the previous SOTA models, RAS shows a 9.7% improvement in short-form question-answering and a 7.9% gain in long-form generation tasks. Additionally, when applied to closed-source Sonnet-3.5 model, RAS consistently achieves superior results compared to single retrieval RAG approaches, including retrieve-and-summarize approach SuRe [25]. We find that sometimes (e.g., on TriviaQA and PubHealth) single-hop retrieval could not boost LLM’s performance, and even makes it worse, which demonstrates the necessity of on-demand retrieval, aligning with previous findings.
Our “role-swapping” study in Figure 3, where Sonnet-3.5 and RAS alternately perform the planning and answering tasks, reveals that performance is primarily bounded by the answering capability. For example, when Sonnet-3.5 handles planning and RAS handles answering, the system achieves 62.4% accuracy on the ARC-Challenge dataset. However, when Sonnet-3.5 is responsible for both planning and answering, accuracy increases to 93.9%. Notably, despite having 60 fewer parameters, our trained RAS achieves planning performance comparable to Sonnet-3.5, and even outperforms it in some cases.
Also, for long-form generation, our trained model achieved notably high MAUVE scores, which can be majorly attributed to our “text-to-triple” step, as demonstrated by our ablation study in Table 2.
4.2 Ablations & More Analysis
We analyze the impact of each component of RAS in Table 2 and provide detailed discussion below.
@@ 4.2 Ablations & More Analysis
| No GraphEncode | 68.8 | 38.7 | 67.3 | 36.5 | 93.6 |
| No Planning | 66.7 | 37.8 | 71.5 | 37.2 | 95.2 |
@@
Effect of LoRA and Multi-task Learning. Our experiments reveal that parameter-efficient training strategies significantly impact model performance. Using only graph token optimization without LoRA leads to substantial degradation (-11.8% on average). A similar observation can be made for “No Multi-Task” (e.g., 12.3% accuracy degradation on PubHealth), indicating the significance of jointly training the model on both action planning and answer generation tasks rather than optimizing for each task separately, supporting findings from prior work [31]. The complementary effects suggest that while graph-based knowledge representation is valuable, it needs to be combined with careful parameter tuning and multi-task learning to achieve optimal performance.
Impact of Graph Information Abundance. Figure 3 shows that increasing the amount of structured graph information consistently improves RAS’s performance across tasks. For TQA, both RAS and RAS exhibit approximately linear gains as more triples are retained, with no clear saturation even at 100% information. For ASQA-MV, RAS benefits from a sharp improvement between 10% and 30% graph information, followed by steady increases, while RAS maintains a smoother and more stable growth pattern throughout. These results confirm that RAS effectively leverages structured knowledge at all levels of availability, and that more complete knowledge graphs consistently translate to stronger reasoning and generation quality. Additionally, the larger 8B model consistently outperforms the 7B variant under all conditions, suggesting that scaling the model size further enhances RAS’s ability to utilize structured knowledge. Interestingly, even partial graphs containing only 30%–50% of triples already deliver substantial gains over low-information baselines, highlighting RAS’s robustness to incomplete or partially retrieved knowledge.
@@
Impact of Training Data Volume. Figure 4 demonstrates how training dataset size influences model performance across different tasks. Considering the computational efficiency, we sampled 2,000 instances each from TQA and 2WQA for evaluation, while maintaining the original sizes of other datasets. The results indicate that model performance generally improves with increased training data volume. Notably, our model achieves competitive performance even with limited training data - using only 5% (10K instances) of our full dataset, it surpasses previous state-of-the-art models on TQA, 2WQA, and ELI5. These results suggest both the robustness of our architectural design and the effectiveness of our data curation methodology. However, we observed an exception with the ELI5 dataset, where performance were inconsistent. This irregularity can be attributed to the inclusion of ASQA training data in our training set, following established setting from previous research [30, 31]. Among our test datasets, ASQA and ELI5 are unique in requiring long-form response generation. The periodic decline in ELI5 performance suggests that the model’s response generation began to align more closely with ASQA’s training data distribution, potentially at the expense of ELI5’s distinct characteristics.
| Triple Extractor | TQA | 2WQA | PopQA |
| Flan-T5-Large | 70.3 | 40.7 | 56.7 |
| LLaMA-3.2-3B | 72.7 | 42.1 | 58.3 |
| Claude-3.5-Sonnet | 73.8 | 44.5 | 60.1 |
Impact of Triple Extraction Quality We further examine how the quality of triple extraction influences the overall effectiveness of RAS. Specifically, we test three text-to-triples models: Flan-T5-Large, LLaMA-3.2-3B model and Claude-3.5-Sonnet where the first two are trained on WikiOfGraph data. As shown in Table 3, higher-quality triple extraction leads to consistent performance gains across all evaluated datasets. Notably, using Claude-3.5-Sonnet triples improves TriviaQA accuracy from 72.7% to 73.8%, boosts 2WikiMultihopQA F1 from 42.1 to 44.5, and raises PopQA accuracy from 58.3% to 60.1%. These results underscore that while RAS achieves strong performance even with lightweight extraction models, its effectiveness can be further enhanced by supplying more precise and complete structured knowledge during inference.